Harry invested a total of $20,000.00 part at 7% and part at 10%. How much did he invest at each rate if the total interest earned in one year was $1640?
Answer by checkley77(12569)

.10x+.07(20,000-x)=1640
.10x+1,400-.07x=1,640
.03x=1,640-1,400
.03x=240
x=240/.03
x=8,000 amount invested @ 10%.
20,000-8,000=12,000 amount invested @ 7%
Proof:
.10*8,000+.07*12,000=1,640
800+840=1,640
1640=1640